OpenStudy (dan815):

Can someone clarify this question... is he investing the 10K as he gets it You have a trust fund that will become available to you three years from now (at the end of the third year). The fund will pay $10,000 every six months starting at the end of the 3rd year for 10 years (to the end of the 13th year). If you invest all of this money into a savings account that has an interest rate of 8% compounded weekly how much money will you have 30 years from now?

OpenStudy (usukidoll):

..........

OpenStudy (dan815):

what -.-

OpenStudy (dan815):

its this random economics course theyre making us all take

OpenStudy (usukidoll):

dang been forever since I solved an economics problem.

OpenStudy (dan815):

dont need to solve it

OpenStudy (dan815):

just need to know what the question is asking, with the assumption i made it became this hugeeee problem

OpenStudy (dan815):

so im rethinking my assumption

ganeshie8 (ganeshie8):

for 10 years, u will be getting 10K 20 times. (twice a year)

ganeshie8 (ganeshie8):

you're investing it to a savings account that pays 8% interest compounded weekly

ganeshie8 (ganeshie8):

for example :- ur first check : 10K at the start of 4th year yields 10K(1+.08/52)^[52*(30-3)] in 30 years from now.

OpenStudy (dan815):

right so

ganeshie8 (ganeshie8):

form a series

OpenStudy (dan815):

everytime he gets the 10k is that added to what he is already investing

ganeshie8 (ganeshie8):

yes, every 6 months he adds 10K to his savings account

OpenStudy (dan815):

alright gotcha thanks

ganeshie8 (ganeshie8):

this is called annuity problem.. simplifies to a nice geometric series

OpenStudy (dan815):

ya i noticedd

OpenStudy (dan815):

i got a formula like this

OpenStudy (dan815):

kj + (kj^2+kj) + (kj^3+kj^2+kj)+... first 2nd 3rd

OpenStudy (the_fizicx99):

*pretends to knows what everyone is say* -nods-

OpenStudy (dan815):

in that form where my j=1+(r/n)^(n*k/20)

OpenStudy (dan815):

does that look right i mean the series

OpenStudy (the_fizicx99):

saying* smh

OpenStudy (dan815):

and i did some random math to try and solve for that series

OpenStudy (dan815):

so there should be 20kj + 19kj^2+18kj^3.... like that, and i solved for a formula that gives me the sum of this

OpenStudy (dan815):

but it was a lot of math, i might have made many random mistakes here and there xD

ganeshie8 (ganeshie8):

4_a : 10K(1+.08/52)^[52*(30-3)] 4_b : 10K(1+.08/52)^[52*(30-3-1/2)] 5_a : 10K(1+.08/52)^[52*(30-3-2/2)] 5_b : 10K(1+.08/52)^[52*(30-3-3/2)] ...

ganeshie8 (ganeshie8):

we need to add them all. does that look right

OpenStudy (dan815):

waiitt how come u can do that

ganeshie8 (ganeshie8):

it simplifies to ur form, im sure..

OpenStudy (dan815):

interestingg i was thinking about this like...

ganeshie8 (ganeshie8):

4_a : im getting 10K in 3years from now. and ima invest it in bank for (30-3) years hmm...

OpenStudy (dan815):

are u allowed to think about it as as 10k being compounded n times + 10K again being compounded n-1 times +10K...

ganeshie8 (ganeshie8):

yes we need to think like that oly. cuz we are getting money every 6 months.

OpenStudy (dan815):

ohh realllyyy

OpenStudy (dan815):

oncee second!! give me 2 mins i need to think about that

ganeshie8 (ganeshie8):

ok.. even i need to think about ur equations.. :)

OpenStudy (dan815):

i dont why but i just kept thinking that cannot be possible.. with the whole compouding thing,, doesnt compounding 20K at once differ from compounding 10k and 10k

ganeshie8 (ganeshie8):

how ? 20K (1 + r/n)^t = 10K(1+r/n)^t + 10K(1+r/n)^t

OpenStudy (dan815):

oh ya that is true but wait thats not what i meant

OpenStudy (dan815):

for the 2nd compound it would be a function of the compound before

OpenStudy (dan815):

so i was thinking how you can write it as separate compounding

OpenStudy (usukidoll):

these business problems are nastay :P

OpenStudy (dan815):

because by the time he is trying to find the compound for the 2nd term, there is some extra money that is being compounded that he isnt concerning himself with right?? or is he calculating that part too with the inital formula

ganeshie8 (ganeshie8):

he is calculating that part too. think of it like this : each 10K packet has its own life. compounds separately itself. you dont need to add every 6 months all ur investments. it all turns out to be same...

ganeshie8 (ganeshie8):

4_a : 10K(1+.08/52)^[52*(30-3)] 4_b : 10K(1+.08/52)^[52*(30-3-1/2)] 5_a : 10K(1+.08/52)^[52*(30-3-2/2)] 5_b : 10K(1+.08/52)^[52*(30-3-3/2)] ...

ganeshie8 (ganeshie8):

4_a : 10K(1+.08/52)^[52*(30-3)] is taking into account compounding effects also for first 10K

OpenStudy (dan815):

ya you are right arrghh

OpenStudy (dan815):

mann i wasted so much time not thinking about this part lol

OpenStudy (dan815):

i did all this crazy math avoiding that simplification

ganeshie8 (ganeshie8):

basically, you can pull out 10K, and wat remains is a simple geometric series..

ganeshie8 (ganeshie8):

lol... i spent more than a week on this a year ago.... to make sense of annuity setup... :o

OpenStudy (dan815):

look at my picture!! its crazy lol i found this really intersting way though

OpenStudy (dan815):

to deal with this sum 20kj+19kj^2+18kj^3....

OpenStudy (dan815):

that will probably come in handy... ahemmm

ganeshie8 (ganeshie8):

im still trying to understand.. lol its mouthful

OpenStudy (dan815):

20kj+19kj^2+18kj^3.... +and- (kj^2+2kj^3....)

OpenStudy (dan815):

so thank i can write it as 20k(j+j^2+j^3...) - k(j^2+2j^3+3j^4....) then repeat the same process on the right side again

OpenStudy (dan815):

then there was this series pattern that emerged out of this lol for the first time i saw the whole + and - alternating series pop out randomly

OpenStudy (dan815):

i found the formula for sum of odd geometric series didnt even think it was possible till i realized this cool simplifcation to be made

OpenStudy (dan815):

k(j^3+j^5+j^7...) =kj(j^2+j^4+j^6...) =kj((j^2)^1+(j^2)^2+(j^2)^3...) = kj(j^2-(J^2)^n/1-j^2)

OpenStudy (dan815):

this is all probably really noob to you though lol

OpenStudy (dan815):

i just found it really fun... to do this question.. i felt like oh wow finally i am using all that random knowledge about gemoetric series haha

ganeshie8 (ganeshie8):

oh you're adding up previous accumulated money to 10K each time, and simplifying it... looks nice :) it comes very handy when u do a problem like below :- Suppose you have taken 500K for loan, and you want to pay back in 10 years at an interest rate of 10% compounded quarterly. Calculate the EMI.

ganeshie8 (ganeshie8):

im going for lunch... cya :)

OpenStudy (dan815):

cya! sleep time for me

ganeshie8 (ganeshie8):

gn have good sleep :D

OpenStudy (dan815):

ohh this is another way... the geometric series formulas will simplify in the end 20k(j+j^2+j^3...) - k(j^2+2j^3+3j^4....) =20k(j+j^2+j^3...) - k(j^2+j^3+j^4....)-k(j^3+j^4+...)-k(j^4+j^5+...)-k(...... =20k(j-j^21/1-j) - k (J^2-j^21/1-j)-k(j^3-j^21/1-j).... =20k(j-j^21/1-j) - [k/(1-j)] *-1*(2+3+4...+20) + 20*(k/1-j)*j^21 =20k([(j-j^21)/(1-j)]- [k/(1-j)]*-1*((22*19)/2) + 20*(k/(1-j))*j^21

OpenStudy (dan815):

well this stuff is pretty much impossible to see without latex.. or drawing